### v1

# BeeManc at the PLABA Track of TAC-2023: Investigating LLMs and Controllable Attributes for Improving Biomedical Text Readability

Abstract
In this system report, we describe the models and methods we used for our participation in the PLABA2023 task on biomedical abstract simplification, part of the TAC 2023 tracks. The system outputs we submitted come from the following three categories: 1) domain fine-tuned T5-like models including Biomedical-T5 and Lay-SciFive; 2) fine-tuned BART-Large model with controllable attributes (via tokens) BART-w-CTs; 3) ChatGPT-prompting. We also present the work we carried out for this task on BioGPT fine-tuning.
In the official automatic evaluation using SARI scores, BeeManc ranks 2nd among all teams and our model Lay-SciFive ranks 3rd among all 13 evaluated systems. In the official human evaluation, our model BART-w-CTs ranks 2nd on Sentence-Simplicity (score 92.84), 3rd on Term-Simplicity (score 82.33) among all 7 evaluated systems; It also produced a high score 91.57 on Fluency in comparison to the highest score 93.53. In the second round of submissions, our team using ChatGPT-prompting ranks the 2nd in several categories including simplified term accuracy score 92.26 and completeness score 96.58, and a very similar score on faithfulness score 95.3 to re-evaluated PLABA-base-1 (95.73) via human evaluations.
Our codes, fine-tuned models, prompts, and data splits from the system development stage will be available at https://github.com/HECTA-UoM/PLABA-MU
1 Introducing the Methodologies
The overall framework of our experimental design is displayed in Figure 1. In the first step, we fine-tune selected Large Language Models (LLMs) including T5, SciFive, BioGPT, and BART, apply prompt-based learning for ChatGPTs, and optimise control mechanisms on BART model. Then, we select the best performing two models using quantitative evaluation metrics SARI, BERTscore, BLEU, and ROUGE. Finally, we chose a subset of the testing results of the two best-performing models for our human evaluation. In this section, we first introduce the models we used, followed by LoRA efficient training, and then introduce the quantitative metrics we applied.
Figure 1: Model Development and Selection. BART is fine-tuned using Wikilarge data. MAX step chooses the two best-performing models according to the automatic evaluation results using SARI and BERTscore.
1.1 Models
The baseline models we investigated in our work include T5, SciFive, GPTs, BioGPT, BART, and Control Mechanisms. We give more details below.
1.1.1 T5
T5 Raffel et al. (2020) used the same Transformer structure from Vaswani et al. (2017) but framed the text-to-text learning tasks using the same vocabulary, sentence piece tokenisation, training, loss, and decoding. The pre-fixed tasks include summarisation, question answering, classification, and translation. T5 used the common crawl corpus and filtered to keep only natural text and de-duplication processing. They extracted 750GB of clean English data to feed into the model for multi-task pre-training. Different masking strategies are integrated into the T5 model to facilitate better performances of specific fine-tuning tasks. Notably, it has been successfully employed in various fields such as Clinical T5 by Lehman and Johnson (2023) and Lu et al. (2022).
Furthermore, T5’s pre-training on an extensive and diverse corpus of text data endows it with a strong foundation in understanding intricate language structures, a crucial asset for handling medical terminology. Its fine-tuning capability further enhances its adaptability, allowing us to tailor the model specifically to the nuances of the biomedical text simplification task. These attributes make T5 a compelling candidate for this work. In this paper, we fine-tuned three versions of T5, namely t5-small, t5-base, and t5-large, paired with their sentence-piece pre-trained tokenizer, which leads to our Biomedical-T5s. Each is fine-tuned independently on the same dataset as the other models to provide comparable results. Note that we use the prompt “summarize:” as it is the closest to our task.
1.1.2 SciFive
Using the framework of T5, SciFive is a Large Language Model pre-trained on the biomedical domain and has demonstrated advanced performances on multiple biomedical NLP tasks Phan et al. (2021).
While preserving the abilities of T5 in sequence-to-sequence tasks, SciFive offers a deep understanding of medical terminology, concepts, and language structures. As a result, SciFive emerges as a strong candidate for text summarisation and medical language processing tasks, offering the potential to generate clear and accurate simplifications of medical texts.
Similarly to our work on T5, we fine-tuned two versions of SciFive, namely SciFive-base and SciFive-large, paired with their pre-trained tokenizer. Each is fine-tuned independently on the same dataset. We again use the prompt “summarize:” for task-relation purposes. Because SciFive itself was already pre-trained in the biomedical domain, we name our fine-tuned SciFive for text simplification task as “Lay-SciFive”.
Figure 2: GPT Prompt Examples
1.1.3 OpenAI’s GPTs
Given the remarkable performance demonstrated by OpenAI’s GPT models in text simplificationJeblick et al. (2022), we decided to apply simplifications using “GPT-3.5-turbo” and GPT-4 via their API [^1] https://openai.com/blog/openai-api. Both models were accessed. Example prompts we used can be found in Figure 2.
1.1.4 BioGPT
BioGPT Luo et al. (2022) is an advanced language model specifically designed for medical text generation. BioGPT is built upon GPT-3 but is specifically trained to understand medical language, terminology, and concepts. BioGPT follows the Transformer language model backbone and is pre-trained on 15 million PubMed abstracts. It has demonstrated a high level of accuracy and has great potential for applications in medicine.
BioGPT is fine-tuned on our training and validation set, as with other encoder-decoder models (Figure 1) producing our Lay-BioGPT.
1.1.5 BART
Like the default transformer structure, BART Lewis et al. (2020) aims to address the issues in BERT and GPT models by integrating their structures with a bi-directional encoder and an autoregressive decoder. In addition to the single token masking strategy applied in BERT, BART provides various masking strategies, including deletion, span masking, permutation, and rotation of sentences. Compared to GPT models, BART provides both leftward and rightward context in the encoders.
1.1.6 Controllable Mechanisms
We applied the modified control token strategy in Li et al. (2022) for both BART-base and BART-large models. The training includes 2 stages, leveraging both Wikilarge training set Zhang and Lapata (2017) and our split of training set from PLABA Attal et al. (2023).
The four attributes for control tokens (CTs) are listed below:
  • <DEPENDENCYTREEDEPTH_x> (DTD)
  • <WORDRANK_x> (WR)
  • <REPLACEONLYLEVENSHTEIN_x> (LV)
  • <LENGTHRATIO_x> (LR)
They represent 1) the syntactic complexity, 2) the lexical complexity, 3) the inverse similarity of input and output at the character level, and 4) the length ratio of input and output respectively.
Before training, the four CTs are calculated and prepared for the 2 stage training sets. In both stages, we pick the best model in 10 epochs based on the training loss of the validation set. We applied the best model from the first stage as the base model and fine-tuned it on our PLABA training set for 10 epochs.
After fine-tuning, the next step is to find the optimal value of CTs. Following a similar process in MUSS Martin et al. (2020), we applied Nevergrad Rapin and Teytaud (2018) on the validation set to find the static optimal discrete value for DTD, WR, and LV. As for the LR, we applied the control token predictor to maximise the performance with the flexible value. The predictor is also trained on Wikilarge data Zhang and Lapata (2017) to predict the potential optimal value for LR. We call the integrated methods of applying CTs on the BART model as BART-w-CTs.
1.2 LoRA and LLMs
To evaluate bigger model architectures, we fine-tune FLAN-T5 XL Chung et al. (2022) and BioGPT-Large, which have 3 billion and 1.5 billion parameters, respectively. FLAN-T5 XL is based on the pre-trained T5 model with instructions for better zero-shot and few-shot performance. To optimise training efficiency, and as our computational resources do not allow us to fine-tune the full version of these models, we employ the LoRA Hu et al. (2022) technique, which allows us to freeze certain parameters, resulting in more efficient fine-tuning with minimal trade-offs.
2 Model Development
The only data set we used for the model fine-tuning development is from PLABA Attal et al. (2023). So, our team belongs to the data-constrained category if other teams applied broader and external biomedical data for their model development.
2.1 Data Preprocessing and Setup
To investigate the selected models for training and fine-tuning, we divided the original PLABA data into Train, Validation, and Test sets, aiming for an 8:1:1 ratio. However, in the real implementations, we found that there are only a few 1-to-0 sentence pairs, which might cause a negative effect in training the simplification models. Thus we eliminated all 1-to-0 sentence pairs. In addition, to better leverage the SARI score, we picked sentences with multi-references for validation and testing purposes. As a result, we ended up with the following sentence pair numbers according to the source sentences (5757, 814, 814). Note that for our submitted three system outputs from BART-w-CTs, Biomedical-T5, and Lay-SciFive, the first one only used our training set split, while the other two used both our training and validation sets. We expect the models will have even better performances if we have re-trained the model using all the PLABA data with the optimised parameters, i.e. including our test set split.
2.2 Automatic Evaluations
In this section, we list quantitative evaluation scores and some explanations for them. The results for T5 Small, T5 Base, T5 Large, FLAN-T5 XL with LORA, SciFive Base, SciFive Large, and BART models with CTs (BART-w-CTS) are displayed in Table 1. Interestingly, the fine-tuned T5 Small model obtains the highest scores in both BLEU and ROUGE metrics including ROUGE-1, ROUGE-2, and ROUGE-L. The fine-tuned BART Large with CTs (BART-L-w-CTs) produces the highest SARI score at 46.54; while the fine-tuned T5 Base model achieved the highest BERTScore (72.62) with a slightly lower SARI score (44.10). The fine-tuned SciFive Large achieved the highest SARI score (44.38) among T5-like models, though it is approximately 2 points lower than BART Large with CTs.
The quantitative evaluation scores of GPT-like models are presented in Table 2 including GPT-3.5 and GPT-4 using prompts, and fine-tuned BioGPT with LoRA. GPT-3.5 reported relatively higher scores than GPT-4 on all lexical metrics except for SARI, and a much higher BERTScore than GPT-4 (58.35 vs 46.99). In comparison, BioGPT-Large with LoRA reported the lowest SARI score (18.44) and the highest BERTScore (62.9) among these three GPT-like models. Comparing the models across Table 1 and Table 2, the GPT-like models did not beat T5-Base on both SARI and BERTScore and did not beat BART-w-CTs on SARI. Therefore, we did not pick GPT-like models for our submission to the shared task, since each team can only submit up to three system outputs.
To look into the details of model comparisons from different epochs on the extracted testing set, we present the learning curve of T5, SciFive, BART-base on WikiLarge (BART-b-Wiki), and continuously-learned BART-b-Wiki on the PLABA data (our training set only) in Figure 3. We also present the learning curve of T5 Base and BART Base using different metrics in Figure 4.
Models BLEU ROUGE-1 ROUGE-2 ROUGE-L SARI BERTScore
  T5 Small 49.86 65.94 48.60 63.94 33.38 69.58
T5 Base 43.92 64.36 46.07 61.63 44.10 72.62
T5 Large 43.52 64.27 46.01 61.53 43.70 60.39
FLAN-T5 XL (LoRA) 44.54 63.16 45.06 60.53 43.47 67.94
SciFive Base 44.91 64.67 46.45 61.89 44.27 60.86
SciFive Large 44.12 64.32 46.21 61.41 44.38 72.59
BART Base with CTs 21.52 56.14 35.22 52.38 46.52 50.53
BART Large with CTs 20.71 54.73 32.64 49.68 46.54 50.16
Table 1: Automatic Evaluations of fine-tuned T5, SciFive, and BART Models with Control Token (BARTs-w-CTs) mechanisms on the Extracted Testing Set. FLAN-T5 XL used LoRA.
Models BLEU ROUGE-1 ROUGE-2 ROUGE-L SARI BERTScore
  GPT-3.5 20.97 50.07 24.72 43.12 42.61 58.35
GPT-4 19.50 48.36 23.34 42.38 43.22 46.99
BioGPT-Large (LoRA) 41.36 63.21 46.63 61.56 18.44 62.9
Table 2: Automatic Evaluations of 1-shot GPTs and fine-tuned BioGPT-Large on the Extracted Testing Set
Figure 3: Learning Curves using BERTScores and SARI Scores of T5, SciFive and BART Models on the Extracted Testing Set
Figure 4: Learning Curves of T5-base and BART-base using BLEU, ROUGE, SARI, and BERTScores on the Extracted Testing Set
Because the fine-tuned T5-Base model has the highest BERTScore (72.62) and also a relatively higher SARI score (44.10), we chose it as one of the candidates for human evaluation. The other candidate is the fine-tuned BART Large with CT mechanisms which has the highest SARI score (46.54) among all evaluated models. Note that, SciFive Large has results close to the ones of T5 Base. In this case, we selected the smaller model for human evaluation.
Original Text Simplified sentences Meaning preservation Simplicity
A national programme of neonatal screening for CAH would be justified, with reassessment after an agreed period. A national program of checking newborns for COVID-19 would be a good idea. Neutral Agree
A national programme of newborn screening for CAH would be justified, with reassessment after an agreed period of time. Strongly Agree Strongly Disagree
Table 3: Sample of human evaluation questionnaire
Model Meaning Preservation Simplicity
BART with CTs 2.625 2.900
T5-base 3.094 2.244
Table 4: Human Evaluation Scores of Fine-tuned BART-Large with CTs and T5-Base Models on the Sampled Test Set
Annotator Meaning preservation Simplicity
0 & 1 0.583 0.138
0 & 2 0.238 0.126
1 & 3 0.008 0.238
2 & 3 -0.130 -0.014
Table 5: Cohen Kappa among annotators over 3 categories ordinal - win, lose, and tie.
Anno. Model Meaning preservation Simplicity
0 & 1 T5-base 0.449 0.386
BART w CTs 0.441 0.052
0 & 2 T5-base 0.259 0.202
BART w CTs 0.200 0.007
1 & 3 T5-base 0.307 0.065
BART w CTs -0.141 -0.056
2 & 3 T5-base -0.056 0.116
BART w CTs 0.065 -0.285
Table 6: Krippendorff’s alpha among annotators (Anno.) over the 5-Likert scale from strongly agree to strongly disagree.
2.3 Human Evaluation
In our human evaluation, we randomly sampled 80 sentences from the test set split and evaluated the corresponding outputs of BART-large with CTs and T5-base in an anonymised setting. In the human evaluation form, we randomly assigned the order of the two systems’ outputs and put them beside the input sentence. Based on the comparison of input and output sentences, the annotators need to answer two questions: “To what extent do you agree the simplified sentence keeps the major information” and “To what extent do you agree the simplified sentence is well simplified”. The answer is limited to a 5-point Likert scale, from strongly disagree to strongly agree. The sample form can be found in Table 3.
There are 4 annotators in this human evaluation, one of the annotators is a native English speaker, and the others are fluent English users as a second language. Two annotators are final-year bachelor students, one annotator is a Masters candidate, and the last one is a Postdoctoral researcher. Each annotator evaluated 40 sentence pairs with 50% overlaps to make sure every sentence was evaluated twice by different annotators. The detailed human evaluation scores on the two selected models we evaluated are shown in Table 4 using the two designed criteria.
Based on the cross-annotation (overlaps), we calculated the inter-rater agreement level in Table 5 using Cohen’s Kappa on cross models and Table 6 using Krippendorff’s Alpha with model-wise comparison. Both tables include the agreement levels on two sub-categories, namely “meaning preservation” and “text simplicity”. Because there is no overlap in the annotation tasks between annotators (0, 3) and annotators (1, 2), we listed the agreement levels between the available pairs of annotations, i.e. (0, 1), (0, 2), (1, 3), and (2, 3). The Cohen’s Kappa agreement level presented in Table 5 shows the inter-rater agreement on the performance order of the two systems, whether one system is better than the other or tie. The Krippendorff’s Alpha represented in Table 6 shows the annotation reliability over all 5-Likert options.
Based on the results from Table 4, annotators show a different preference for the two systems. Despite the limited gap in the SARI score, BART with CTs shows a better capability to conduct the simplification tasks. Yet regarding meaning preservation, the fine-tuned T5-base performs better, as the BERTScore tells among the comparisons in Table 1.
From Table 5, Annotators 0 and 1 have the highest agreement on “meaning preservation” (score 0.583), while Annotators 1 and 3 have the highest agreement on “simplicity” (score 0.238) evaluation across the two models. This also indicates that it is not easy to evaluate the system performances against each other regarding these two criteria.
Model wise, Table 6 further shows that Annotators 0 and 1 agree more on the judgements of fine-tuned T5 Base model on both two criteria of “meaning preservation” (score 0.449) and “text simplicity” (score 0.386), in comparison to the BART model. On the fine-tuned BART Large model with CTs, these two annotators only agreed on the “meaning preservation” factor with a score of 0.441. This phenomenon also applies to Annotators 0 and 2 regarding the judgement of these two models. Interestingly, while Table 5 shows that Annotators 1 and 3 have a better agreement on “simplicity” judgement over “meaning preservation” using Cohen’s Kappa, Table 6 shows the opposite using Krippendorff’s Alpha, i.e. it tells the agreement on “meaning preservation” of these two annotators instead. This shows the difference between the two agreement and reliability measurement metrics.
3 BeeManc Submissions to PLABA2023
3.1 1st-round submissions
For system submissions, we chose our domain and task fine-tuned BART-w-CTs as our first system, Biomedical-T5, and Lay-SciFive as our second and third ones.
From the official human evaluation in Figure 5 (upper part), our submission BART-w-CTs ranks 2nd on Simp.sent 92.84, 3rd on simp.term 82.33, among the 7 evaluated systems. Note that BART-w-CTs also produced very comparable scores with the highest system 91.57 vs 93.53, even though it ranked 5th on simp.fluency.
From the official automatic evaluations using 4-reference-based SARI scores in Figure 6, BeeManc team ranked the 2nd after valeknappich team among all teams. Our model BeeManc -3 (Lay-SciFive) ranks 3rd among all system submissions with a score 0.420299 after the systems valeknappich-1/2. Note that our other two systems BeeManc -1/2 (BART-w-CTs and Biomedical-T5) also achieved 0.40+ scores that are above all other systems behind us.
Figure 5: The official Human Evaluation on Simplicity, Accuracy, Fluency, Completeness, and Faithfulness at sentence-level, term-level, and averages. PLABA-base-1 was re-evaluated in round-2 submissions. Ondov et al. (2024)
Figure 6: Official SARI Scores from Teams computed against 4 references (submission round 1). BeeManc 1/2/3: BART-w-CTs, Biomedical-T5, Lay-SciFive Ondov et al. (2024)
3.2 2nd-round submissions
Due to the availability of human evaluators, there was a second call for the alternative system selection for human evaluation purposes by the task organisers. To testify the GPT-like models, we submitted our 4th system outputs using the ChatGPT prompting, which produced a system-level automatic evaluation SARI score 0.37. However, looking at the human evaluation outcomes in Figure 5 (lower part), ChatGPT-prompting wins the second highest score on several categories including Simp.term.acc score 92.26 and Acc. comp. score 96.58, and a very similar score on Acc. faith. score 95.3 to re-evaluated PLABA-base-1 (95.73). The drawbacks of ChatGPT-prompting are the sentence and term simplification, with under 0.80 scores. Overall, ChatGPT-prompting generates better accuracy, fluency, and faithfulness scores but falls on sentence and term simplifications.
4 Conclusions and Future Work
In this system description, we have reported our model development using LLMs and Control Mechanisms for the text simplification task on biomedical abstract using the PLABA data set and our submissions to PLABA-2023 track. From three of our submitted systems BART-w-CTs, Biomedical-T5, and Lay-SciFive, we have the following highlights in the official outcomes:
  • Our team (BeeManc ) ranks 2nd among all teams in the automatic evaluation using the SARI score.
  • Our system Lay-SciFive ranks 3rd out of 13 evaluated systems using the SARI score.
  • BART-w-CTs produced 2nd and 3rd highest scores on sentence-simplicity and term-simplicity.
  • Our second round submission ChatGPT-prompting achieved the 2nd highest score on several categories including Simp.term.acc and Acc. comp.
In future work, we plan to carry out investigations on more recent models including BioBART Yuan et al. (2022), try different prompting methods such as the work from Cui et al. (2023), and design a more detailed human evaluation such as the work proposed by Gladkoff and Han (2022) with error severity levels might shed some light on this.
For this shared task challenge, we carried out data-constrained fine-tuning and model development only using the PLABA data set. However, in the broader sense, we would like to carry out data-augmented training such as using biomedical synthetic data generated by state-of-the-art Transformer-like models Belkadi et al. (2023).
Author Contributions
ZL carried out experiments on BART and Control Mechanisms, prepared Human Evaluation data, Measured inter-rater agreement. SB carried out experiments on T5 and SciFive. NM carried out experiments on GPTs and Bio-GPTs. LH supervised the project and helped drafting the system report. Human evaluation by ZL, SB, NM, and LH. Writing-wise, Abstract: LH; Section 1: ZL, SB, NM, LH; Section 2: ZL, SB, LH; Section 3: LH; Section 4: LH. Appendix: SB and LH. MS and GN: co-supervisors and approved the manuscript.
Acknowledgements
We thank XC for the valuable discussions, thank Brian Ondov and Hoa T. Dang for keeping in touch with us during the shared task. LH and GN are grateful for the support from the grant “Assembling the Data Jigsaw: Powering Robust Research on the Causes, Determinants and Outcomes of MSK Disease”. The project has been funded by the Nuffield Foundation, but the views expressed are those of the authors and not necessarily the Foundation. Visit www.nuffieldfoundation.org. LH and GN were also supported by the grant “Integrating hospital outpatient letters into the healthcare data space” (EP/V047949/1; funder: UKRI/EPSRC).